OpenStudy (anonymous):

Solve the given system, using the substitution method. y = 3x 7 6x 2y = 12 (Points : 1) a.) (14, 12) b.) (12, 14) c.)There is no solution. d.)There are an infinite number of solutions.

OpenStudy (mathmate):

The sign in the second equations are not clear (i.e. not visible). They make a big difference. You can get the answer by substituting x in each equation and see if you get the right value of y.

OpenStudy (anonymous):

6x-2y =12

OpenStudy (mathmate):

Sorry, don't see the sign of the first either.

OpenStudy (anonymous):

it is minus as well: y=3x-7

OpenStudy (mathmate):

y=3x-7 => 3x-y=7 => 6x-2y = 14 Compare with 6x-2y=12 what do you see, and what do you conclude?

OpenStudy (anonymous):

No Algebraic Solution solution 1. The value of a variable that makes an equation true. 2. In chemistry, a solution is a homogeneous mixture composed of only one phase.

OpenStudy (mathmate):

We're maths here, so no (algebraic) solution is correct.